please help me solve this equation(4r-2)-(3r=1)
Answer by checkley77(12844)

(4r-2)-(3r=1)
4r-2-3r=-1
r-2=-1
r=-1+2
r=1 answer.
UNLESS you meant:
(4r-2)-(3r)=1 then:
4r-2-3r=1
r=1+2
r=3 answer.